OpenStudy (anonymous):

12x = 24 what is x ?

OpenStudy (anonymous):

24/12

OpenStudy (anonymous):

x=24/12

OpenStudy (shubhamsrg):

x= 132/66 ! ;)

OpenStudy (anonymous):

lol

mathslover (mathslover):

\[\huge{12x=24}\] \[\huge{\frac{12x}{12}=\frac{24}{12}}\] \[\huge{x=\frac{24}{12}}\] \[\huge{x=2}\]